# Rethinking Debiasing: Real-World Bias Analysis and Mitigation

### v3

@@ E.3 Application of DiD on the bias detection task
Table 9: We experiment with three feature destruction methods with various hyper-parameters on HMLP distributed dataset with LfF.
$T_{fd}$ param BC Avg
N/A N/A 47.70 ± 3.58 54.15 ± 3.02
pixel-shuffle 1 51.44 ± 1.01 55.43 ± 0.20
patch-shuffle 2 51.07 ± 0.48 55.29 ± 0.27
4 49.41 ± 0.26 55.40 ± 0.26
8 54.81 ± 0.74 63.06 ± 0.77
16 49.74 ± 1.10 53.69 ± 0.31
center-occlusion 8 45.19 ± 1.41 51.61 ± 1.31
16 47.26 ± 0.54 50.94 ± 0.59
24 49.00 ± 0.80 52.60 ± 0.55
32 52.44 ± 0.87 55.76 ± 0.16
E.4 Results of BN samples under LMLP settings
@@
E.5 Addtional results on the WaterBirds dataset
@@ Appendix F Related Works
Generally, spurious correlation refers to the phenomenon that two distinct concepts are statistically correlated within the training distribution, though there is no causal relationship between them, e.g. background and foreground object [chu2024task]. The spurious correlation is a vital aspect of understanding how machine learning models learn and generalize [48]. Specifically, studies on distribution shift [6] claim that spurious correlation is one of the major types of distribution shift in the real world, and thus an important distribution shift that a reliable model should be robust to. Furthermore, studies on fairness and bias [53] have demonstrated the pernicious impact of spurious correlation in classification [36], conversation [54], and image captioning [23]. However, despite its broad impact, spurious correlation is generally used as a vague concept in previous works and lacks a proper definition and deeper understanding of it. This is also the major motivation of this work.
@@
In this work, we focus only on debiasing methods that do not require bias information, i.e. without annotation on the spurious attribute, for it is more practical. Existing work [12, 9, 15, 55, 17, 18, 20, 56] in the area generally involve a biased auxiliary model to capture biases within the training data, according to which the debiased is trained with various techniques. We call such paradigm debiasing with biased auxiliary model (DBAM). Specifically, Nam et al. [12] is the first work that follows the DBAM paradigm, proposing to use GCE for bias capture, and the loss-based sample re-weighing scheme to train the debiased model. Lee et al. [9] further proposed a feature augmentation technique to further utilize the captured bias, enhancing the BC samples. Hwang et al. [55] proposed to augment biased data identified according to the biased auxiliary model by applying mixup [57] to contradicting pairs. Lim et al. [17] proposed to conduct adversarial attacks on the biased auxiliary model to augment BC samples aiming to increase the diversity of BC samples. Lee et al. [20] proposed to first filter out BC samples before training the biased auxiliary model aiming to enhance the bias capture process of the biased model. Liu et al. [16] regard the samples misclassified by the biased auxiliary model as BC samples and emphasize them during training of the debiased model. Recently, Park et al. [56] proposed to provide models with explicit spatial guidance that indicates the region of intrinsic features according to a biased auxiliary model. Kim et al. [14] create images without bias attributes using an image-to-image translation model [58] built upon a biased auxiliary model. A recent pair-wise debiasing method $\chi^{2}$ model [59] based on biased auxiliary models encourages the debiased model to retain intra-class compactness using samples generated via feature-level interpolation between BC and BA samples.
@@ Appendix G Limitations and Future Work
We uncover the insufficiency of existing debiasing benchmarks theoretically and empirically, highlighting the importance of debiasing on real-world biases. We further proposed a feature-destruction-based method that focuses on DBAM methods. However, there are still a few limitations of this work: As shown in section E, while our proposed approach effectively improves the performance of existing DBAM methods on all biased distributions from the real world, the performance is still far from satisfactory, which remains to be further improved in future works. We see potential within those limitations and leave them for future research.
@@ Appendix H Boarder Impact
From a technical standpoint, our research provides a comprehensive framework for analyzing and mitigating biases in datasets. The proposed fine-grained analysis framework and evaluation benchmarks offer a new perspective on how biases manifest in real-world data and how existing debiasing methods can be improved. Our approach, which involves the destruction of target features during bias capture, demonstrates significant improvements in handling real-world biases, as evidenced by our extensive experimental results.
By advancing the understanding of dataset biases and improving the performance of debiasing methods, our research contributes to the development of more robust and generalizable AI models. This is particularly relevant in an era where AI systems are increasingly deployed in dynamic and diverse environments, necessitating models that can adapt and maintain high performance across different contexts and populations.

### v4

@@ E.5 Application of DiD on the bias detection task
Table 11: Existing debiasing methods perform poorly on unbiased training data, while DiD greatly boosts the performance.
Algorithm Colored MNIST Corrupted CIFAR10
ERM 94.14 ± 0.21 67.91 ± 0.13
LfF 70.19 ± 1.50 52.04 ± 2.14
+ DiD 93.18 ± 0.26 57.29 ± 0.22
DisEnt 75.24 ± 3.40 58.50 ± 0.20
+ DiD 92.24 ± 0.44 64.58 ± 0.02
BEL 84.14 ± 0.61 55.64 ± 0.66
+ DiD 90.02 ± 0.54 56.28 ± 0.45
BED 80.66 ± 0.90 58.57 ± 0.12
+ DiD 89.10 ± 1.28 62.97 ± 0.16
E.6 Results of BN samples under LMLP settings
@@
E.7 Addtional results on the WaterBirds dataset
@@ Appendix F Related Works
Generally, spurious correlation refers to the phenomenon that two distinct concepts are statistically correlated within the training distribution, though there is no causal relationship between them, e.g. background and foreground object. The spurious correlation is a vital aspect of understanding how machine learning models learn and generalize [52]. Specifically, studies on distribution shift [6] claim that spurious correlation is one of the major types of distribution shift in the real world, and thus an important distribution shift that a reliable model should be robust to. Furthermore, studies on fairness and bias [57] have demonstrated the pernicious impact of spurious correlation in classification [40], conversation [58], and image captioning [26]. However, despite its broad impact, spurious correlation is generally used as a vague concept in previous works and lacks a proper definition and deeper understanding of it. This is also the major motivation of this work.
@@
In this work, we focus only on debiasing methods that do not require bias information, i.e. without annotation on the spurious attribute, for it is more practical. Existing work [12, 9, 15, 59, 17, 18, 23, 60, 20, 19] in the area generally involve a biased auxiliary model to capture biases within the training data, according to which the debiased is trained with various techniques. Specifically, Nam et al. [12] is the first work to propose to use GCE for bias capture, and the loss-based sample re-weighing scheme to train the debiased model. Lee et al. [9] further proposed a feature augmentation technique to further utilize the captured bias, enhancing the BC samples. Hwang et al. [59] proposed to augment biased data identified according to the biased auxiliary model by applying mixup [61] to contradicting pairs. Lim et al. [17] proposed to conduct adversarial attacks on the biased auxiliary model to augment BC samples aiming to increase the diversity of BC samples. Lee et al. [23] proposed to first filter out BC samples before training the biased auxiliary model aiming to enhance the bias capture process of the biased model. Liu et al. [16] regards the samples misclassified by the biased auxiliary model as BC samples and emphasizes them during training of the debiased model. Park et al. [60] proposed to provide models with explicit spatial guidance that indicates the region of intrinsic features according to a biased auxiliary model. Kim et al. [14] create images without bias attributes using an image-to-image translation model [62] built upon a biased auxiliary model. A recent pair-wise debiasing method $\chi^{2}$ model [63] based on biased auxiliary models encourages the debiased model to retain intra-class compactness using samples generated via feature-level interpolation between BC and BA samples. Recently, Han et al. [20] propose to use the disagreement leverages the disagreement probability between the target label and the prediction of a biased model to determine the weight of each training example. Sreelatha et al. [19] propose a strategy where tehy train a deep debiased model utilizing the information acquired from both deep (perfectly biased) and shallow (weak debiased) network in the previous phase. Despite different technique routes, it’s implicitly or explicitly assumed by these works that the biases can be well captured by the biased models, which serves as a foundation for subsequent debiasing. However, in this work, we show that such assumptions might be challenged under real-world scenarios.
@@ Appendix G Limitations and Future Work
We uncover the insufficiency of existing debiasing benchmarks theoretically and empirically, highlighting the importance and novel challenges of debiasing in the real world, i.e. Sparse bias capturing. We further proposed a simple yet effective method to address the challenge. However, there are still a few limitations of this work:
  • While we have proposed fine-grained empirical and theoretical analysis on real-world biases with important characteristics found, due to the complexity of the data bias problem, there might be other important characteristics of real-world biases that we are unaware of. We believe this is another important direction for future research, which serves as the foundation for developing debiasing methods that are applicable in the real world.
  • While DiD has been shown to be simple and effective, it remains a preliminary solution to tackle the Sparse bias capturing challenge in real-world debiasing, and there is still much room for improvement. We believe there will be more sophisticated and potentially better-performing approaches to tackle the challenge in the future.
We see potential within those limitations and leave them for future research.
@@ Appendix H Boarder Impact
From a technical standpoint, our research provides a fine-grained framework for analyzing data biases, a systematic evaluation framework for real-world debiasing, and a simple yet effective solution to the challenges in real-world debiasing. The bias analysis framework serves as the basis for deepen our understanding of dataset biases. The evaluation framework paves the path towards developing debiasing methods applicable in real-world scenarios. The proposed approach DiD, is highly effective and adaptive to the various debiasing methods. Thus, we believe DiD have high potential to be adopted in debiasing methods in the future.
By advancing the understanding of dataset biases and improving the performance of debiasing methods in real-world scenarios, our research contributes to the development of more robust and generalizable AI models. This is particularly relevant in an era where AI systems are increasingly deployed in dynamic and diverse environments, necessitating models that can adapt and maintain high performance across different contexts and populations.
Table 13: Experiments on multiple bias setting with LMLP and HMLP combined has demonstrated consistent effectiveness of our approach in handling multiple biases.
Algorithm LMLP BC HMLP BC Avg
LfF 47.09 48.56 48.27
+ DiD +9.40 +10.66 +9.74
DisEnt 50.82 49.15 51.09
+ DiD +6.70 +7.01 +7.00
BEL 53.66 56.33 54.72
+ DiD +2.23 +2.89 +2.69
BED 56.26 54.66 56.54
+ DiD +1.02 +1.20 +0.94
Table 14: Our approach still effectively boosts the performance of even very recent methods. This further demonstrated the adaptability of our approach. The experiments are conducted based on Corrupted CIFAR10.
LMLP HMLP HMHP
Algorithm BC acc Avg acc BC acc Avg acc BC acc Avg acc
DPR 51.54 54.25 43.67 44.67 25.92 31.77
+ DiD +6.97 +4.36 +5.44 +13.16 +2.61 +2.47